a meter stick is held perpendicular to the ground. it forms a shadow that is 1.8 m long. at the same time, a flagpole forms a shadow that is 7.2 m long. how tall is the flagpole?
Answer by rfer(16322)

1/1.8=x/7.2
1.8x=7.2
x=7.2/1.8
x=4 m